# DuoCast: Duo-Probabilistic Diffusion for Precipitation Nowcasting

### v3

@@ Abstract
Accurate short-term precipitation forecasting is critical for weather-sensitive decision-making in agriculture, transportation, and disaster response. Existing deep learning approaches often struggle to balance global structural consistency with local detail preservation, especially under complex meteorological conditions. We propose DuoCast, a dual-diffusion framework that decomposes precipitation forecasting into low- and high-frequency components modeled in orthogonal latent subspaces. We theoretically prove that this frequency decomposition reduces prediction error compared to conventional single branch U-Net diffusion models. In DuoCast, the low-frequency model captures large-scale trends via convolutional encoders conditioned on weather front dynamics, while the high-frequency model refines fine-scale variability using a self-attention-based architecture. Experiments on four benchmark radar datasets show that DuoCast consistently outperforms state-of-the-art baselines, achieving superior accuracy in both spatial detail and temporal evolution.
@@ Introduction
Precipitation plays a vital role in weather systems, influencing temperature, humidity, and overall atmospheric conditions. Precipitation nowcasting, which predicts short-term rainfall using recent radar observations, is increasingly important for applications such as agriculture, transportation, and disaster response (Zhang et al. 2023). Conventional approaches to nowcasting rely on numerical weather prediction (NWP), which simulates atmospheric dynamics by solving partial differential equations (PDEs) derived from physical laws (Bauer, Thorpe, and Brunet 2015; Lorenc 1986; Alley, Emanuel, and Zhang 2019). However, the complexity of NWP renders them computationally expensive and inefficient for real-time forecasting (Bi et al. 2023).
Figure 1: Challenges in precipitation nowcasting: a) precipitation with weather (warm) front patterns, and b) high-frequency features (micro-scale variability) within edge regions shown in red parallelograms.
Recent advances in data-driven deep learning techniques have shown greater potential than traditional NWP methods in precipitation prediction by leveraging large datasets and avoiding the need to solve complex physical equations. Deterministic models (Shi et al. 2017; Gao et al. 2022b; Yan et al. 2024; Lin et al. 2025) aim to capture the overall motion of precipitation systems by generating mean-value forecasts. However, they often produce blurry predictions and lack fine-grained spatial detail (Ravuri et al. 2021). In contrast, probabilistic models (Ravuri et al. 2021; Zhang et al. 2023; Gao et al. 2024) introduce stochastic latent variables to better capture local-scale variability. While effective at modeling uncertainty, treating the entire system as fully stochastic can introduce excessive randomness and degrade forecasting precision (Yu et al. 2024). To address this, hybrid approaches such as DiffCast (Yu et al. 2024) combine global deterministic forecasting with local probabilistic refinement. Nevertheless, its CNN-based diffusion model struggles to preserve high-frequency information due to limited capacity (Yang et al. 2023), resulting in less spatially accurate precipitation intensity, as illustrated in Fig. 1(b).
@@
Precipitation nowcasting has traditionally relied on numerical weather prediction (NWP), which solves atmospheric PDEs (Molteni et al. 1996; Skamarock et al. 2008). Although effective, NWP is computationally intensive, often requiring hours on supercomputers (Bi et al. 2023). Recent deep learning models such as LDCast (Leinonen et al. 2023) and MetNet-3 (Andrychowicz et al. 2023) offer competitive performance with significantly lower latency. They are generally categorized as either deterministic or probabilistic.
Deterministic Precipitation Models
aim to predict the mean-based evolution of precipitation by modeling spatial and temporal dynamics. ConvLSTM (Shi et al. 2015) captures these dynamics via convolutional and recurrent structures, while PhyDNet (Guen and Thome 2020) introduces a physics-guided decomposition into stochastic and deterministic motion. FACL (Yan et al. 2024) replaces standard L2 loss with a loss tailored for signal-based spatiotemporal modeling. AlphaPre (Lin et al. 2025) disentangles phase and amplitude to separately model motion and intensity.
@@
Frequency-domain analysis has also shown strong effectiveness in the image analysis and generation domain (Tan et al. 2024; Woo et al. 2022; Wen et al. 2023; Chen et al. 2025), serving as a valuable complement to spatial-based approaches, which often struggle with specific artifacts. For instance, (Doloriel and Cheung 2024) apply frequency-based masking to extract shared features, while (Zhou et al. 2024) categorize frequency components into semantic, structural, and noise levels to identify regions with unique spectral characteristics.
Figure 2: Overview of the proposed DuoCast framework. DuoCast models both low- and high-frequency precipitation processes by leveraging meteorological signals—particularly weather fronts—from historical observations. It consists of two diffusion stages: a low-frequency process of convolutions that captures front-guided precipitation trends, and a high-frequency process of self-attentions that refines micro-scale variability for precise, pixel-level forecasting.
@@
To enhance precipitation estimation, DuoCast employs two distinct stochastic diffusion processes to separately model low- and high-frequency components in dedicated subspaces. An overview of the overall architecture is presented in Fig. 2. Specifically, we design two complementary diffusion models: low-frequency, which integrates weather front information to model low-frequency, large-scale precipitation patterns; and high-frequency, which captures high-frequency, localized variability to refine fine-grained precipitation details.
@@
Low-frequency components in precipitation maps capture the overall scale and structure of precipitation systems, as macro trends. To estimate $\mathbf{Y_{\text{low}}}$ in the low-frequency subspace, a low-frequency diffusion network $\epsilon_{\theta_{\text{low}}}$ is employed, where $\theta_{\text{low}}$ indicates the parameters of the diffusion model. For forward diffusion, we sample from $q(\mathbf{Y_{\text{low}}^{\text{t}}}|\mathbf{Y_{\text{low}}^{\text{0}}})$ in a closed form at an arbitrary timestep $t\in T$ as $q(\mathbf{Y_{\text{low}}^{\text{t}}}|\mathbf{Y_{\text{low}}^{\text{t-1}}})=\mathcal{N}(\mathbf{Y_{\text{low}}^{\text{t}}};\sqrt{1-\beta_{t}}\mathbf{Y_{\text{low}}^{\text{t-1}}},\beta_{t}\mathbf{I})$. The reverse diffusion process is formulated as:
@@
High-frequency components correspond to localized precipitation intensities, shown as micro-scale variations. To estimate $\mathbf{Y}_{\text{high}}$, we introduce a high-frequency diffusion model $\epsilon_{\theta_{\text{high}}}$, which follows a similar formulation as the low-frequency counterpart, as defined in Equations 3 and 5. To optimize ${\epsilon}_{\theta_{\text{high}}}$, an objective function is defined as follows:
@@
While convolutional backbones like U-Net are commonly used in diffusion models, they are limited in generating high-frequency components—particularly crucial in precipitation nowcasting. We first illustrate the limitations of existing architectures in modeling high-frequency patterns, and then explain how a two-stage design operating in distinct spectral subspaces addresses this issue. A detailed version of the analysis, including formal proofs, is provided in the supplementary material.

### v4

@@ Abstract
Accurate short-term precipitation forecasting is critical for weather-sensitive decision-making in agriculture, transportation, and disaster response. Existing deep learning approaches often struggle to balance global structural consistency with local detail preservation, especially under complex meteorological conditions. We propose DuoCast, a dual-diffusion framework that decomposes precipitation forecasting into low- and high-frequency components modeled in orthogonal latent subspaces. We theoretically prove that this frequency decomposition reduces prediction error compared to conventional single branch U-Net diffusion models. In DuoCast, the low-frequency model captures large-scale trends via convolutional encoders conditioned on weather front dynamics, while the high-frequency model refines fine-scale variability using a self-attention-based architecture. Experiments on four benchmark radar datasets show that DuoCast outperforms state-of-the-art baselines, achieving superior accuracy in both spatial detail and temporal evolution.
Codehttps://github.com/ph-w2000/DuoCast
Appendixhttps://arxiv.org/abs/2412.01091
@@ Introduction
Precipitation plays a vital role in weather systems, influencing temperature, humidity, and overall atmospheric conditions. Precipitation nowcasting, which predicts short-term rainfall using recent radar observations, is increasingly important for applications such as agriculture, transportation, and disaster response (Zhang et al. 2023). Its importance was underscored during the July 2025 Texas flash floods: National Weather Service (NWS) observed localized totals over 5 inches and rainfall rates of 2–3 inches per hour at many locations, yet forecasts slightly mispredicted the heaviest cells, resulting in flash flooding in central Texas and trapping hundreds of people (Moore et al. 2025). Conventional approaches to nowcasting rely on numerical weather prediction (NWP), which simulates atmospheric dynamics by solving partial differential equations (PDEs) derived from physical laws (Bauer et al. 2015; Lorenc 1986; Alley et al. 2019). However, the complexity of NWP renders them computationally expensive and inefficient for real-time forecasting (Bi et al. 2023).
Figure 1: Challenges in precipitation nowcasting: a) precipitation with weather (warm) front patterns, and b) high-frequency features (micro-scale variability) within edge regions shown in pink parallelograms.
Recent advances in data-driven deep learning techniques have shown greater potential than traditional NWP methods in precipitation prediction by leveraging large datasets and avoiding the need to solve complex physical equations. Deterministic models (Shi et al. 2017; Gao et al. 2022; Yan et al. 2024; Lin et al. 2025) aim to capture the overall motion of precipitation systems by generating mean-value forecasts. However, they often produce blurry predictions and lack fine-grained spatial detail (Ravuri et al. 2021). In contrast, probabilistic models (Ravuri et al. 2021; Zhang et al. 2023; Gao et al. 2024) introduce stochastic latent variables to better capture local-scale variability. While effective at modeling uncertainty, treating the entire system as fully stochastic can introduce excessive randomness and degrade forecasting precision (Yu et al. 2024). To address this, hybrid approaches such as DiffCast (Yu et al. 2024) combine global deterministic forecasting with local probabilistic refinement. Nevertheless, its CNN-based diffusion model struggles to preserve high-frequency information due to limited capacity (Yang et al. 2023), resulting in less spatially accurate precipitation intensity, as illustrated in Fig. 1(b).
@@
Precipitation nowcasting has traditionally relied on numerical weather prediction (NWP), which solves atmospheric PDEs (Skamarock et al. 2008). Although effective, NWP is computationally intensive, often requiring hours on supercomputers (Bi et al. 2023). Recent deep learning models (Andrychowicz et al. 2023) offer competitive performance with significantly lower latency. They are generally categorized as either deterministic or probabilistic.
Deterministic Precipitation Models aim to predict the mean-based evolution of precipitation by modeling spatial and temporal dynamics. ConvLSTM (Shi et al. 2015) captures these dynamics via convolutional and recurrent structures, while PhyDNet (Guen and Thome 2020) introduces a physics-guided decomposition into stochastic and deterministic motion. FACL (Yan et al. 2024) replaces standard L2 loss with a loss tailored for signal-based spatiotemporal modeling. AlphaPre (Lin et al. 2025) disentangles phase and amplitude to separately model motion and intensity.
@@
Frequency-domain analysis has shown strong effectiveness in image analysis and generation (Tan et al. 2024; Woo and others 2022; Wen et al. 2023; Chen et al. 2025), serving as a valuable complement to spatial-based approaches, which often struggle with artifacts. For instance, (Doloriel and Cheung 2024) apply frequency-based masking to extract shared features, while (Zhou et al. 2024) categorize frequency components into semantic, structural, and noise levels to identify regions with unique spectral characteristics.
Figure 2: Overview of DuoCast. A two-stage diffusion model leveraging historical weather fronts signals: (i) a low-frequency convolutional stage capturing front-guided precipitation trends, and (ii) a high-frequency self-attention stage refining micro-scale variability for pixel-level forecasts.
@@
To enhance precipitation estimation, DuoCast employs two distinct stochastic diffusion processes to model low- and high-frequency components in dedicated subspaces. The architecture overview is presented in Fig. 2. Specifically, we design two complementary diffusion models: low-frequency, which integrates weather front information to model low-frequency, large-scale precipitation patterns; and high-frequency, which captures high-frequency, localized variability to refine fine-grained precipitation details.
@@
Low-frequency components capture the overall scale and structure of precipitation systems, as macro trends. To estimate $\mathbf{Y_{\text{low}}}$ in the low-frequency subspace, a low-frequency diffusion network $\epsilon_{\theta_{\text{low}}}$ is employed, where $\theta_{\text{low}}$ indicates the parameters of the diffusion model. For forward diffusion, we sample from $q(\mathbf{Y_{\text{low}}^{\text{t}}}|\mathbf{Y_{\text{low}}^{\text{0}}})$ in a closed form at an arbitrary timestep $t\in T$ as $q(\mathbf{Y_{\text{low}}^{\text{t}}}|\mathbf{Y_{\text{low}}^{\text{t-1}}})=\mathcal{N}(\mathbf{Y_{\text{low}}^{\text{t}}};\sqrt{1-\beta_{t}}\mathbf{Y_{\text{low}}^{\text{t-1}}},\beta_{t}\mathbf{I})$. The reverse diffusion process is formulated as:
@@
High-frequency components correspond to localized precipitation intensities, shown as micro-scale variations. To estimate $\mathbf{Y}_{\text{high}}$, we introduce a high-frequency diffusion model $\epsilon_{\theta_{\text{high}}}$, which follows a similar formulation as the low-frequency counterpart, as defined in Equations 3 and 4. To optimize ${\epsilon}_{\theta_{\text{high}}}$, an objective function is defined as follows:
@@
While convolutional backbones like U-Net are commonly used in diffusion models, they are limited in generating high-frequency components—particularly crucial in precipitation nowcasting. We first illustrate the limitations of existing architectures in modeling high-frequency patterns, and then explain how a two-stage design operating in distinct spectral subspaces addresses this issue.
$\bigl|\widehat{k}(\xi)\bigr|\;\leq\;{C_{\mathrm{BV}}}/({1+|\xi|}),\qquad\forall\,\xi\in\mathbb{R}^{d}.$
$H_{L}(\xi):=\prod_{\ell=1}^{L}\widehat{k_{\ell}}(\xi).$ (6)
Theorem 1 (Spectral envelope under bounded variation).
Under the above assumptions, there exists $C>0$ such that
Sharp Tail Bound.
We now derive an input–dependent high–frequency estimate that retains the true spectrum.
Theorem 2 (Sharp capacity bottleneck).
Corollary 1 (Irreducible high‑frequency error).
so the approximation error cannot be eliminated by any depth‑$L$ CNN.
Two–Stage Diffusion Modeling.
For any target $\mathbf{Y}=\mathbf{Y}_{\text{low}}+\mathbf{Y}_{\text{high}}$, where $\mathbf{Y}_{\text{low}}\in\mathcal{B}_{\omega_{c}}$ and $\mathbf{Y}_{\text{high}}\in\mathcal{B}_{\omega_{c}}^{\perp}$, assume the first and second diffusion model generation $g^{(1)}_{\theta}\colon\mathcal{B}_{\omega_{c}}\!\to\!\mathcal{B}_{\omega_{c}}$ and $g^{(2)}_{\phi}\colon\mathcal{B}_{\omega_{c}}^{\perp}\!\to\!\mathcal{B}_{\omega_{c}}^{\perp}$, respectively. Orthogonality yields:
Corollary 2 (Two–stage universal approximation).
We next investigate two (approximated) architectural designs for the low- and high-frequency diffusion models.
Low-Frequency Diffusion Model
Weather fronts, which evolve gradually and span broad spatial regions, are key to shaping precipitation systems (Catto et al. 2012). For instance, warm fronts bring steady light rain, cold fronts induce abrupt heavy rainfall, and occluded fronts result in widespread precipitation. Inadequate modeling of them can lead to significant forecasting errors. As shown in Fig. 1(a), the context indicates a weakening warm front, yet DiffCast erroneously predicts intensifying rainfall. Therefore, the low-frequency diffusion captures these trends with convolutions to guide prediction, conditioned on front dynamics derived from underlying air mass interactions.
Air Mass Spatial Modeling.
To model the spatial organization of weather fronts, we introduce a convolutional module that learns the distribution and interactions of underlying air masses. As air mass boundaries often coincide with sharp gradients in precipitation (Lagerquist et al. 2019), this module extracts spatial cues relevant to front structure. Specifically, we apply three tailored convolutional layers to the preliminary forecasts $\bar{\mathbf{Y}}$ and $\mathbf{Y}^{\dagger}$ to capture local context, multi-scale structure, and temporal anchoring, forming a spatial representation $\mathbf{A}$.
$\mathbf{A}=\text{Concat}(\bar{\mathbf{A}},\mathbf{A}^{\dagger}).$
Weather Front Temporal Modeling.
[…]
High-Frequency Diffusion Model
Optimization
$\mathcal{L}=\lambda_{1}\mathcal{L}_{P}+\lambda_{2}\mathcal{L}_{\text{low}}+\lambda_{3}\mathcal{L}_{\text{high}},$ (10)
where $\lambda_{1}$, $\lambda_{2}$ and $\lambda_{3}$ are hyperparameters.
Experiments & Discussions
Experimental Settings
Datasets
To evaluate DuoCast’s effectiveness in generating precise precipitation maps, we performed experiments using four radar echo datasets: SEVIR (Veillette et al. 2020), MeteoNet (Larvor and Berthomier 2021), Shanghai_Radar (Chen et al. 2020), and CIKM [^1] https://tianchi.aliyun.com/dataset/1085.
Method SEVIR
↑CSI-M ↑CSI-181 ↑CSI-219 ↑HSS ↑SSIM
DuoCast 0.338 0.182 0.107 0.432 0.683
High-Freq 0.310 0.158 0.050 0.401 0.611
Low-Freq 0.327 0.164 0.066 0.410 0.642
Low-Frequency Backbone Exploration
Attention 0.310 0.143 0.043 0.393 0.608
Conv 0.327 0.164 0.066 0.410 0.642
- $\text{Conv}_{\text{I}}$ 0.316 0.155 0.051 0.405 0.631
- $\text{Conv}_{\text{II}}$ 0.317 0.159 0.048 0.400 0.632
- $\text{Conv}_{\text{III}}$ 0.320 0.156 0.050 0.404 0.622
- $\mathbf{Y^{\dagger}}$ 0.316 0.146 0.048 0.400 0.619
- Airmass 0.309 0.143 0.051 0.388 0.625
- Front 0.311 0.138 0.050 0.383 0.628
Evaluation Metrics
Following SOTA (Yu et al. 2024; Lin et al. 2025; Gong et al. 2024), we evaluate the nowcasting with the average Critical Success Index (CSI), Heidke Skill Score (HSS), Structural Similarity Index Measure (SSIM). The CSI, similar to the Intersection over Union (IoU), measures the degree of pixel-wise alignment between predictions and ground truth after thresholding them into binary (0/1) matrices, while HSS reflects improvement over random guessing. Following (Lin et al. 2025), we measure and average CSI (CSI-M) and CSI with two highest thresholds to measure the performance of high-intensity precipitation forecasting. SSIM assesses prediction quality.
Implementation Details
DuoCast was trained with Adam optimizer (learning rate $1\times 10^{-4}$) and 1,000 diffusion steps per low-/high-frequency branch. We used a two-stage schedule: first train the low-frequency model, then jointly train both. The learning rate was fixed across stages; to stabilize training we applied loss weights. In the first stage, the loss weight coefficients were $\lambda_{1}=\lambda_{2}=0.5$. In the second stage, the loss weight were $\lambda_{1}=\lambda_{2}=0.1$ and $\lambda_{3}=0.5$. The SimVP model is used as initial predictor $P$. The cutoff $\theta_{\text{init}}$ is set as the high-intensity precipitation threshold defined in the datasets.
Compared with the State-of-the-Art
To assess the performance of our DuoCast framework in precipitation nowcasting, we compare it against a range of state-of-the-art methods, including probabilistic, deterministic, and hybrid approaches.
Method SEVIR MeteoNet
↑CSI-M ↑CSI-181 ↑CSI-219 ↑HSS ↑SSIM ↑CSI-M ↑CSI-24 ↑CSI-32 ↑HSS ↑SSIM
MAU (2021) 0.3076 0.1071 0.0516 0.3863 0.6505 0.3233 0.2839 0.0997 0.4452 0.7897
SimVP (2022) 0.3108 0.1106 0.0517 0.3924 0.6508 0.3351 0.3002 0.1130 0.4573 0.7804
FourCastNet (2022) 0.2686 0.0717 0.0339 0.3355 0.5976 0.3027 0.2533 0.1085 0.4216 0.6450
Earthformer (2022) 0.2892 0.0844 0.0245 0.3665 0.6633 0.3205 0.2884 0.1237 0.4491 0.7772
PhyDNet (2020) 0.3017 0.1040 0.0278 0.3812 0.6532 0.3384 0.3194 0.1366 0.4673 0.7823
EarthFarseer (2024) 0.3004 0.0992 0.0413 0.3829 0.6327 0.3404 0.3170 0.1372 0.4726 0.7542
NowcastNet (2023) 0.2791 0.0770 0.0351 0.3512 0.6839 0.3427 0.3206 0.1598 0.4751 0.7879
PreDiff (2024) 0.2744 0.0627 0.0235 0.3592 0.5881 0.2944 0.2641 0.1202 0.4462 0.7085
DiffCast (2024) 0.3050 0.1300 0.0582 0.3996 0.6482 0.3512 0.3340 0.1808 0.4846 0.7887
CasCast (2024) 0.2878 0.0954 0.0312 0.3563 0.5680 0.3317 0.3021 0.1351 0.4682 0.7622
FACL (2024) 0.3161 0.1349 0.0655 0.4033 0.5686 0.3612 0.3341 0.1801 0.5006 0.7881
AlphaPre (2025) 0.3259 0.1332 0.0545 0.4110 0.6884 0.3824 0.3633 0.2002 0.5164 0.7968
Ours 0.3375 0.1818 0.1074 0.4318 0.6827 0.3892 0.3841 0.2381 0.5297 0.7981
Method Shanghai Radar CIKM
↑CSI-M ↑CSI-35 ↑CSI-40 ↑HSS ↑SSIM ↑CSI-M ↑CSI-35 ↑CSI-40 ↑HSS ↑SSIM
MAU (2021) 0.3983 0.3621 0.2417 0.5346 0.7195 0.3039 0.2054 0.1241 0.3928 0.6325
SimVP (2022) 0.3850 0.3549 0.2382 0.5194 0.7795 0.3052 0.2044 0.1321 0.3955 0.6538
FourCastNet (2022) 0.3571 0.3108 0.2073 0.4868 0.5598 0.2980 0.1849 0.1015 0.3801 0.4359
Earthformer (2022) 0.3503 0.3178 0.1872 0.4844 0.7298 0.3077 0.2039 0.1369 0.4001 0.6267
PhyDNet (2020) 0.3654 0.3236 0.2176 0.4957 0.7751 0.3038 0.2052 0.1287 0.3931 0.6541
EarthFarseer (2024) 0.3926 0.3608 0.2343 0.5330 0.5405 0.3000 0.2046 0.1259 0.3911 0.6373
NowcastNet (2023) 0.3953 0.3608 0.2450 0.5334 0.7902 0.2991 0.1940 0.1188 0.3865 0.6713
PreDiff (2024) 0.3504 0.3026 0.1855 0.4991 0.6981 0.2842 0.1848 0.1002 0.3481 0.5582
DiffCast (2024) 0.4089 0.3740 0.2606 0.5476 0.7879 0.3159 0.2009 0.1457 0.4085 0.6499
CasCast (2024) 0.3651 0.3291 0.2192 0.4971 0.7250 0.3021 0.1974 0.1142 0.3806 0.5241
FACL (2024) 0.3940 0.3511 0.2355 0.5308 0.7601 0.3105 0.1993 0.1354 0.3996 0.6442
AlphaPre (2025) 0.4178 0.3854 0.2615 0.5534 0.7951 0.3194 0.2068 0.1416 0.4137 0.6568
Ours 0.4252 0.3933 0.2821 0.5643 0.7788 0.3179 0.2275 0.1562 0.4217 0.6659
Quantitative Analysis
Qualitative Analysis
Figure 3: Qualitative comparison with SoTA on a SEVIR event. DuoCast captures finer micro-scale details (pink box) and preserves more consistent evolution trends (pink line).
Ablation Study
Low-Frequency Diffusion Model.
High-Frequency Diffusion Model.
Air Mass Modeling.
As shown in Table 1, removing air mass spatial modeling leads to a notable drop in CSI, particularly CSI-M, indicating its importance in capturing the spatial evolution of weather fronts. Additionally, we analyze the effectiveness of convolution in air mass spatial modeling. As shown in the Fig. 5, $\text{Conv}_{\text{I}}$, captures the spatial dynamics of surrounding air mass effects, highlighted by the heatmap. The frontal convolution, $\text{Conv}_{\text{II}}$, focuses on occluded fronts, which occur in critical transition areas where cold and warm air masses interact, as depicted by the figure’s red band. The temporal convection convolution, $\text{Conv}_{\text{III}}$, captures pixel-level temporal convection, with the red areas indicating potential temporal motion. Quantitatively, as in Table 1, every component in air mass spatial modeling contributes to superior performance.
Weather Front Modeling.
$\mathbf{Y^{\dagger}}$ for High-Intensity Precipitation
To assess the role of $\mathbf{Y^{\dagger}}$ in modeling high-intensity precipitation, we conduct targeted experiments on SEVIR. As shown in Table 1, removing $\mathbf{Y^{\dagger}}$ results in substantial performance drops for severe rainfall events, with CSI-181 and CSI-219 decreasing by 10.8% and 27.0%, respectively. Figure 4(d) visualizes high-intensity events (values $\geq$ 181), where predictions with $\mathbf{Y^{\dagger}}$ align closely with the ground truth, consistently maintaining the precipitation band over time. In contrast, models without $\mathbf{Y^{\dagger}}$ increasingly miss key regions, and the predicted band degrades notably after 30 minutes. These results highlight the importance of $\mathbf{Y^{\dagger}}$ in learning temporal consistency and structure in severe precipitation. Figure 6(b) shows consistently higher CSI across thresholds, confirming our model’s strength in forecasting high-intensity rainfall.
Figure 6: CSI across different (a) lead time steps and (b) intensity thresholds for different methods on SEVIR.
Conclusion
We introduce DuoCast—a dual-probabilistic meteorology-aware model designed for preserving different frequency information. Low-frequency aims to capture the broader trend of weather patterns and high-frequency is to obtain high frequency micro-scale variability relevant to precipitation. Comprehensive experiments on four real-world datasets validate the effectiveness of the proposed framework.
Acknowledgments
This work was supported by the Australian Research Council (ARC) Linkage Project #LP230100294 and ECU Science Early Career and New Staff Grant Scheme.
Supplementary Material
Appendix A Motivating Observation
Precipitation nowcasting signals contain a high proportion of high-frequency components, which capture micro-scale intensity variations but are difficult for conventional convolution-based diffusion models to preserve. To address this, we propose DuoCast, a dual-diffusion framework that decomposes the forecasting task into separate low- and high-frequency components, each modeled within orthogonal latent subspaces for frequency-specific representation and generation. We quantify the frequency characteristics of precipitation data by applying the Discrete Fourier Transform (DFT) and following the approach of (Yang et al. 2025), where high-frequency components are defined using a threshold parameter $\mu$. As shown in Table 3, the SEVIR dataset contains nearly twice the proportion of high-frequency content compared to ImageNet when $\mu=0.5$, highlighting the importance of modeling high-frequency signals in this domain.
$\mu$ $0.5$
ImageNet 0.005561
CIFAR100 0.005200
MSCOCO 0.005760
SEVIR 0.010059
Table 3: The proportion of high-frequency components in different dataset.
Appendix B Formal Proofs
This section presents the detailed theoretical analysis and complete proofs of the main results in the manuscript.
Foundations of Spectral Subspace Diffusion
Spectral Decay of Finite Support Convolutions.
While convolutional backbones like U-Net are commonly used in diffusion models, they are limited in generating high-frequency components—particularly crucial in precipitation nowcasting. We first illustrate how these architectures fall short in modeling such fine-scale patterns.